From GlueXWiki
Jump to: navigation, search

October 10, 2013 FDC meeting


  1. Installation preparations [1] (all)
  2. Engineering (Bill)
  3. Electronics (Chris, Nick)
  4. Tests in EEL126
  5. Other


Participants: Eugene, Bill,Dave, Chris, Simon, Fernando, Beni, and Lubomir.

Installation preparations

- Pictures linked above show the current status of the FDC platform in the Hall. The first cart has been installed on the platform and the survey people aligned the rails inside BCAL with the rails on this cart. The cooling manifold for the BCAL on the upstream side of the magnet has not been installed yet but should be done soon (Bill). The racks on the left side (looking downstream) have been put in place, probably not finally installed. The holes below the racks on the right side have been marked but not cut. The mesh has not been modified yet. Holes in the beam below the platform have been made on the left side (right will be done soon, as well) and we discussed if the size is big enough for one bundle of 8/9 cables; the problem is that the connectors at this side are big and you have to un-bundle the end to insert the cables one by one. We will have ~18 bundles at one side, not sure if we have 18 holes. Also, we will need more people when pulling/pushing the cables through these holes.

- We discussed what to do with the systems that are absolutely needed for the FDC installation but will not be ready before the installation starts: gas, electric power and cooling. Gas: Dave will check what has been installed and how it can be modified. Beni: if some of the 500 ccpm flow controllers have been installed already then we don't have enough for a temporary gas system. Electric power: for the LV crate we need 208V which would require a portable transformer. Fernando volunteered to check the availability and what else is required to power the LV crate (the minimum needed to start cabling).

- The moving date will depend on the readiness of the above (temporary or permanent) systems. In any case it can't be next week (Eugene) because in case of a shutdown the FDC better be in the EEL. Eugene needs information from Lubomir and Beni to include the FDC check (of the gas flow) in the lists of the critical tasks in case of a shutdown. The AC in 126 is already included.


- The skin between the FDC packages has been installed by Bill and Casey, also Tina today. Tomorrow (probably) Casey will install the tubes between the cooling manifolds.


- Half of the signal cables have been labeled by the summer students and we decided to label the rest. The cables for one package were moved from Blue Crab in August and can be labeled. The rest of the cables (also one package) has been used for the full electronics test. We started disconnecting these cables and most are packed back in their original boxes in 126. When Beni is done with the F1TDC test next week we will finish the rest.

Tests at EEL126

- Beni showed his results of the F1TDC tests (Entry 117 in the Elog) in the following configuration: two regular cells (3 and 6) with +HV of 2150V, but (due to the lack of HV cables) only half of each cell was connected to negative -500V, the other half put on 0V (ground). One can clearly see the difference in the drift time spectra (last two plots in the entry). Some of the cards/ASICs (8 chan.) are noisy, and 3 channels are missing most likely due to the cards (we have used the last spares). Fernando suggested to try using the pulser built in the new F1TDCs, first, to test this feature, and second, to identify the source of the missing channels. The same pulser is in fADC125 but has not been tested (firmware not ready yet).

- Eugene started discussion about the cathode deformation during the tensioning and the effects on the resolution. Beni explained that he sees some deviations from the uniformity in terms of the wire positions as reconstructed from the strips. Eugene has measured the elongation perpendicular to the strips to be 0.25% and much bigger than along the strips (GlueX doc 1090), but for 5micron copper. Based on his estimations in the same doc (Table 2) for 2 micron copper the elongation must be factor of 2 bigger but very similar in both directions, i.e. about 5mm in total. This certainly has to be taken into account in the analyses.